In a Nitrogen metabolism study, the following data were obtained over the test period:

Nitrogen consumed 175 grams

Total Fecal N 35 grams

Metabolic Fecal N 10 grams

Total Urinary N 80 grams

Endogenous Urinary N 10 grams

Calculate a. Apparent Nitrogen digestibility (%)

b. True Nitrogen digestibility (%)

c. Nitrogen Balance (g)

d. Apparent Nitrogen Retention (%)

e. Apparent Biological Value (%)

f. True Biological Value (%)

In the example above, the nitrogen retained is calculated in item c. If the gain the animal makes is composed of 20% protein, then how many grams and how many pounds did the animal gain in this time peroid? (there are 454 grams in a pound). Hint: If all of the gain was due to protein, then you would have calculated the gain in c (times 6.25) above. Some amount of gain that is 20% protein is necessary to give you the protein gained, derived from c above.

_____ grams _____ pounds
Answer #1

Answer:

a. Apparent Nitrogen digestibility (%):

(Nitrogen consumed - total fecal nitrogen) / nitrogen consumed x 100%

= (175 - 35) / 175 x 100% = 80%

b.True Nitrogen digestibility (%):

(Nitrogen consumed - (total fecal nitrogen - metabolic fecalnitrogen)) / Nitrogen consumed x 100%

= (175 - (35 - 10)) / 175 x 100% = 85.71%

c. Nitrogen Balance (g):

(Nitrogen consumed - Total fecal nitrogen - total urinary nitrogen)

= 175 - 35 - 80 = 60 g

d. Apparent Nitrogen Retention (%):

(Nitrogen consumed - Total fecal nitrogen - total urinarynitrogen) / Nitrogen consumed x 100%

= (175 - 35 - 80) / 175 x 100% = 34.29%

e. Apparent Biological Value (%):

(Nitrogen consumed - total fecal nitrogen - total urinary nitrogen) / (Nitrogen consumed – total fecal nitrogen) x 100%

= (175 - 35 - 80) / (175 - 35) x 100% = (60 / 140) x 100% = 42.85%

f. True Biological Value (%):

(Nitrogen consumed - (total fecal nitrogen - metabolic fecal nitrogen) - (urinary nitrogen - endogenous urinary nitrogen)) / (Nitrogen consumed – (total fecal nitrogen - metabolic fecal nitrogen) x 100%

= (175 - (35 - 10) - (80 - 10)) / (175 - (35 - 10)) x 100%

= (175 - 25 - 70) / (175 - 25) x 100% = 53.3%

If the gain the animal makes is composed of 20% protein, then the number of grams and amount of pounds the animal gained in this time peroid:

60g x 6.25 x 20% = 60 x 6.25 x 0.2 = 75 grams

75 / 454= 0.17 lb (pounds)